OpenStudy (anonymous):

Can someone explain to me why (2x-2)/5 times 10 equals 2(2x-2)?

hartnn (hartnn):

\(\huge \frac{2x-2}{5}\times 10=(2x-2)\frac{10}{5}\) now what is 10/5= ?

OpenStudy (anonymous):

2! Thank you!!!

hartnn (hartnn):

welcome ^_^